# The Uniform Mordell–Lang Conjecture

### v3

@@ Appendix A Rémond’s theorem revisited
(7.2) #$\#X^{\circ}(\overline{\mathbb{Q}})\cap\Gamma\leq c(g,\deg_{L}X,\deg_{L}A)^{\mathrm{rk}\Gamma+1}.$
Moreover, we may and do assume that
(Property c): As a function, $c$ is increasing in all three invariables.
@@ A.1. Preliminary setup
It follows from Tate’s construction that there exists a constant $c_{\mathrm{NT}}\geq 0$, which depends on $A$, such that $|\hat{h}_{L}(P)-h(P)|\leq c_{\mathrm{NT}}$ for all $P\in A(\overline{\mathbb{Q}})$.
Let $h_{1}$ denote the Weil height of the polynomials defining the addition and the substraction on $A$.
It is known that
(A.1) $c_{\mathrm{NT}},h_{1}\leq c^{\prime}(g,\deg_{L}A)\max\{1,h_{\mathrm{Fal}}(A)\}.$
See [DP07, equation (6.41)]. Alternatively this can be deduced from [DGH21, (8.4) and (8.7)].
@@ A.3. Generalized Mumford’s Inequality
Denote by 00 the origin of the abelian variety $A$. We may and do assume that the stabilizer of $X$ in $A$, denoted by $\mathrm{Stab}(X)$ has dimension 00; otherwise $X^{\circ}=\emptyset$ and the proposition trivially holds true.
@@
Let $C_{1},\ldots,C_{s}$ be the irreducible components of $W$ passing through 00 with $\dim C_{j}=\dim X-i_{0}\geq 1$. Then $s\leq\sum_{j=1}^{s}\deg C_{j}\leq(\deg X)^{i_{0}+1}\leq(\deg X)^{\dim X}$ by Bézout’s Theorem. Moreover,
@@
So (A.3) is equivalent to $\Xi\subseteq\bigcup_{j=1}^{s}\bigcap_{c\in C_{j}(\overline{\mathbb{Q}})}(X-c)$. Each $\bigcap_{c\in C_{j}(\overline{\mathbb{Q}})}(X-c)$ is a finite intersection of at most $\dim X$ members because of dimension reasons. So $\deg\bigcap_{c\in C_{j}(\overline{\mathbb{Q}})}(X-c)\leq(\deg X)^{\dim X}$ by Bézout’s Theorem. Moreover, each irreducible component of $\bigcap_{c\in C_{j}(\overline{\mathbb{Q}})}(X-c)$ has dimension $<\dim X$ because $\dim\mathrm{Stab}(X)=0$. Hence we are in case (ii) by setting $X^{\prime}=\bigcup_{j=1}^{s}\bigcap_{c\in C_{j}}(X-c)$. So we are done. ∎
@@ A.4. Removing $h(X)$
We start by reducing to the case where
(Hyp): $X$ generates $A$.
Indeed, let $A^{\prime}$ be the abelian subvariety of $A$ generated by $X-X$. Then $X\subseteq A^{\prime}+Q$ for some $Q\in A(\overline{\mathbb{Q}})$. The subgroup $\Gamma^{\prime}$ of $A(\overline{\mathbb{Q}})$ generated by $\Gamma$ and $Q$ has rank $\leq\mathrm{rk}\Gamma+1$. We have $(X-Q)^{\circ}=X^{\circ}-Q$ by definition of the Ueno locus, $(X^{\circ}(\overline{\mathbb{Q}})-Q)\cap\Gamma\subseteq(X^{\circ}(\overline{\mathbb{Q}})-Q)\cap\Gamma^{\prime}=X^{\circ}(\overline{\mathbb{Q}})\cap\Gamma^{\prime}$ and $\deg_{L}(X-Q)=\deg_{L}X$. By Lemma 2.5, (Hyp) yields $\deg_{L}A^{\prime}\leq c^{\prime}(g,\deg_{L}X)$. Therefore if (7.2) holds true for $X-Q\subseteq A^{\prime}$, $L|_{A^{\prime}}$ and $\Gamma^{\prime}\cap A^{\prime}(\overline{\mathbb{Q}})$, then #$\#X^{\circ}(\overline{\mathbb{Q}})\cap\Gamma\leq c(g,\deg_{L}X)^{\mathrm{rk}\Gamma^{\prime}+1}\leq c(g,\deg_{L}X)^{\mathrm{rk}\Gamma+2}$. So we can conclude by replacing $c$ with $c^{2}$. Thus we are reduced to the case where (Hyp) holds true.
Now, assume (Hyp). We prove (7.2) by induction on
@@
For an arbitrary $r\geq 1$. Assume (7.2) holds true for $10,1,\ldots,r-$ 1.
@@
Step 1 Handle large points by both inequalities of Rémond. The goal of this step is to prove the following bound: there exists a constant $c_{6}=c_{6}(g,\deg_{L}X,\deg_{L}A)>0$ such that
(A.5) #$\#\left\{P\in X^{\circ}(\overline{\mathbb{Q}})\cap\Gamma:\hat{h}_{L}(P)\geq c_{5}\max\{1,h(X),h_{\mathrm{Fal}}(A)\}\right\}\leq c_{6}^{\mathrm{rk}\Gamma+1}.$
@@
Consider the subset $\Xi_{j}=\{P_{j+1},\ldots,P_{j+N^{\prime}}\}$ with $j\in\{0,\ldots,N-N^{\prime}\}$; it has $N^{\prime}$ pairwise distinct elements. We claim that $\Xi$ cannot be contained in a proper Zariski closed subset $X^{\prime}\subsetneq X$ with $\deg_{L}X^{\prime}\leq(\deg_{L}X)^{2\dim X}$. Indeed if such an $X^{\prime}$ exists, then by definition of the Ueno locus we have $(X^{\prime})^{\circ}\supseteq X^{\circ}\cap X^{\prime}$. So $\Xi\subseteq(X^{\prime})^{\circ}(\overline{\mathbb{Q}})\cap\Gamma$. As $\dim X^{\prime}<\dim X=r$, we can apply the induction hypothesis (7.2) to each irreducible component of $X^{\prime}$. As $X^{\prime}$ has $\leq(\deg_{L}X)^{2r}$ irreducible components and each component has degree $\leq(\deg_{L}X)^{2r}$, we then get #$\#\Xi\leq(\deg_{L}X)^{2r}c(g,(\deg_{L}X)^{2r},\deg_{L}A)^{\mathrm{rk}\Gamma+1}$. This contradicts our choice of $N^{\prime}$ because $c$ is increasing in all the three variables (Property c).
@@
Recall that we have covered $\Gamma\otimes\mathbb{R}$ by at most $\lfloor(1+({8c_{4}})^{1/2})^{\mathrm{rk}\Gamma}\rfloor$ cones and each cone contains $<gMN^{\prime}$ points $P\in X^{\circ}(\overline{\mathbb{Q}})\cap\Gamma$ with $\hat{h}_{L}(P)=|P|^{2}\geq c_{5}\max\{1,h(X),h_{\mathrm{Fal}}(A)\}$. Thus
#$\#\left\{P\in X^{\circ}(\overline{\mathbb{Q}})\cap\Gamma:\hat{h}_{L}(P)\geq c_{5}\max\{1,h(X),h_{\mathrm{Fal}}(A)\}\right\}$ $\leq(1+({8c_{4}})^{1/2})^{\mathrm{rk}\Gamma}gMN^{\prime}$
@@
Step 2 Remove the dependence on $h(X)$. More precisely, set
@@
(A.10) #$\#\left\{P\in X^{\circ}(\overline{\mathbb{Q}})\cap\Gamma:\hat{h}_{L}(P-P_{0})\geq c_{8}^{2}\max_{1\leq i\leq N^{\prime\prime}}\hat{h}_{L}(P_{i}-P_{0})+c_{9}\max\{1,h_{\mathrm{Fal}}(A)\}\right\}\leq c_{10}^{\mathrm{rk}\Gamma+1}.$
The proof follows closely [Rém00a, Prop.3.6]. We wish to apply Lemma A.4 to $X-P_{0}$ and the set $S=\{P_{i}-P_{0}:0\leq i\leq N^{\prime\prime}\}$. Let us verify the hypothesis. Let $Y\subseteq X-P_{0}$ with $Y$ equidimensional of dimension $\dim X-1=r-1$ and $S\subseteq Y(\overline{\mathbb{Q}})$, and set $Y^{\prime}:=Y+P_{0}$. Then $P_{i}\in(Y^{\prime})^{\circ}(\overline{\mathbb{Q}})\cap\Gamma$.[6][6][6]As $Y^{\prime}\subseteq X$, we have $(Y^{\prime})^{\circ}\supseteq X^{\circ}\cap Y^{\prime}$ by definition of the Ueno locus. So $P_{i}\in(Y^{\prime}\cap X^{\circ})(\overline{\mathbb{Q}})\cap\Gamma\subseteq(Y^{\prime})^{\circ}(\overline{\mathbb{Q}})\cap\Gamma$. Each irreducible component of $Y^{\prime}$ has dimension $\leq r-1$. Thus we can apply induction hypothesis (7.2) to each irreducible component of $Y^{\prime}$. So $N^{\prime\prime}\leq\sum_{Y^{\prime\prime}}c(g,\deg_{L}Y^{\prime\prime},\deg_{L}A)^{\mathrm{rk}\Gamma+1}$, with $Y^{\prime\prime}$ running over all irreducible components of $Y^{\prime}$. Since $\deg_{L}Y=\deg_{L}Y^{\prime}=\sum_{Y^{\prime\prime}}\deg_{L}Y^{\prime\prime}$ and $c$ is increasing in all the three variables (Property c), this bound implies
@@
We claim $\deg_{L}Y>\deg_{L}A(\deg_{L}X)^{2}/g!$. Indeed, assume $\deg_{L}Y\leq\deg_{L}A(\deg_{L}X)^{2}/g!$. Then $N^{\prime\prime}\leq\deg_{L}A(\deg_{L}X)^{2}/g!\cdot c(g,\deg_{L}A(\deg_{L}X)^{2}/g!,\deg_{L}A)^{\mathrm{rk}\Gamma+1}$ because $c$ is increasing in all the three variables (Property c). This contradicts the definition of $N^{\prime\prime}$ from (A.9).
@@
Step 3 Prove the following alternative.
@@
    Either #$\#X^{\circ}(\overline{\mathbb{Q}})\cap\Gamma\leq N^{\prime\prime}(8c_{8}+1)^{\mathrm{rk}\Gamma}+c_{10}^{\mathrm{rk}\Gamma+1}$;
@@
    or there exists $Q\in X^{\circ}(\overline{\mathbb{Q}})\cap\Gamma$ such that #$\#\{P\in X^{\circ}(\overline{\mathbb{Q}})\cap\Gamma:\hat{h}_{L}(P-Q)\geq 2c_{9}\max\{1,h_{\mathrm{Fal}}(A)\}\}\leq c_{10}^{\mathrm{rk}\Gamma+1}$.
The proof follows closely [Rém00a, Prop.3.7]. Assume we are not in case (i), i.e. #$\#X^{\circ}(\overline{\mathbb{Q}})\cap\Gamma>N^{\prime\prime}(8c_{8}+1)^{\mathrm{rk}\Gamma}+c_{10}^{\mathrm{rk}\Gamma+1}$. Let $c_{12}$ be the smallest real number such that there exists $Q\in X^{\circ}(\overline{\mathbb{Q}})\cap\Gamma$ with
(A.11) #$\#\{P\in X^{\circ}(\overline{\mathbb{Q}})\cap\Gamma:|P-Q|\geq c_{12}\}\leq c_{10}^{\mathrm{rk}\Gamma+1}.$
Consider the set $\Xi:=\{P\in X^{\circ}(\overline{\mathbb{Q}})\cap\Gamma:|P-Q|\leq c_{12}\}$ in the $\mathrm{rk}\Gamma$-dimensional Euclidean space $(\Gamma\otimes\mathbb{R},|\cdot|)$. Then #$\#\Xi\geq\#X^{\circ}(\overline{\mathbb{Q}})\cap\Gamma-\#\{P\in X^{\circ}(\overline{\mathbb{Q}})\cap\Gamma:|P-Q|\geq c_{12}\}>N^{\prime\prime}(4c_{8}+1)^{\mathrm{rk}\Gamma}$. By an elementary ball packing argument, $\Xi$ (being a subset of $\Gamma\otimes\mathbb{R}$ contained in a closed ball of radius $c_{12}$ centered at $Q$) is covered by at most $(8c_{8}+1)^{\mathrm{rk}\Gamma}$ balls of radius $c_{12}/4c_{8}$ centered at points in $X^{\circ}(\overline{\mathbb{Q}})\cap\Gamma$; see [Rém00a, Lem.6.1]. By the Pigeonhole Principle, one of the balls contains $\geq N^{\prime\prime}+1$ points in $X^{\circ}(\overline{\mathbb{Q}})\cap\Gamma$, say $P_{0},P_{1},\ldots,P_{N^{\prime\prime}}$. We have $|P_{i}-P_{0}|\leq c_{12}/2c_{8}$ for each $i$. Thus (A.10) yields
#$\#\left\{P\in X^{\circ}(\overline{\mathbb{Q}})\cap\Gamma:\hat{h}_{L}(P-P_{0})\geq c_{12}^{2}/4+c_{9}\max\{1,h_{\mathrm{Fal}}(A)\}\right\}\leq c_{10}^{\mathrm{rk}\Gamma+1}.$
@@
Step 4 Conclude by the standard packing argument.
Recall our assumption (Hyp) that $X$ generates $A$. The assumption of Theorem 7.1 says that (Hyp pack) holds true for $X$ and $\Gamma$, i.e. we have (7.1).
Assume we are in case (i) from Step 3. Recall the definition of $N^{\prime\prime}=c_{7}^{\mathrm{rk}\Gamma+1}+1$ from (A.9). Then #$\#X^{\circ}(\overline{\mathbb{Q}})\cap\Gamma\leq c^{\mathrm{rk}\Gamma+1}$ for $c:=2\max\{(c_{7}+1)(8c_{8}+1),c_{10}\}$. Hence we can conclude for this case.
@@
(A.12) #$\#\left\{P\in X^{\circ}(\overline{\mathbb{Q}})\cap\Gamma:\hat{h}_{L}(P-Q)\leq 2c_{9}\max\{1,h_{\mathrm{Fal}}(A)\}\right\}\leq c_{0}(1+2\sqrt{2c_{9}c_{0}})^{\mathrm{rk}\Gamma}.$
Thus we have #$\#X^{\circ}(\overline{\mathbb{Q}})\cap\Gamma\leq c_{0}(1+2\sqrt{2c_{9}c_{0}})^{\mathrm{rk}\Gamma}+c_{10}^{\mathrm{rk}\Gamma+1}$. So #$\#X^{\circ}(\overline{\mathbb{Q}})\cap\Gamma\leq c^{\mathrm{rk}\Gamma+1}$ for $c:=2\max\{c_{0},1+2\sqrt{2c_{9}c_{0}},c_{10}\}$. Hence we can conclude are this case.

### v4

@@ Appendix A Rémond’s theorem revisited
(7.2) $X^{\circ}(\overline{\mathbb{Q}})\cap\Gamma\leq c(g,\deg_{L}X,\deg_{L}A)^{\mathrm{rk}\Gamma+1}.$
Moreover, we may and do assume that the function $c$ is increasing in all three invariables, by replacing $c(g,d,l)$ by $\max_{g^{\prime}\leq g,d^{\prime}\leq d,l^{\prime}\leq l}c(g^{\prime},d^{\prime},l^{\prime})$.
@@ A.1. Preliminary setup
Let $c_{\mathrm{NT}}(A)$ and $h_{1}(A)$ be as from [DP07, Thm. 6.8]. It is known that
(A.1) $c_{\mathrm{NT}}(A),h_{1}(A)\leq c^{\prime}\max\{1,h_{\mathrm{Fal}}(A)\}$
for some $c^{\prime}=c^{\prime}(g,\deg_{L}A)>0$; see [DP07, equation (6.41)].
@@ A.3. Generalized Mumford’s Inequality
Denote by 0 the origin of the abelian variety $A$. We may and do assume that the stabilizer of $X$ in $A$, denoted by $\mathrm{Stab}(X)$ has dimension 0; otherwise $X^{\circ}=\emptyset$ and the proposition trivially holds true.
@@
Let $C_{1},\ldots,C_{s}$ be the irreducible components of $W$ passing through 0 with $\dim C_{j}=\dim X-i_{0}\geq 1$. Then $s\leq\sum_{j=1}^{s}\deg C_{j}\leq(\deg X)^{i_{0}+1}\leq(\deg X)^{\dim X}$ by Bézout’s Theorem. Moreover,
@@
So (A.3) is equivalent to $\Xi\subseteq\bigcup_{j=1}^{s}\bigcap_{c\in C_{j}(\overline{\mathbb{Q}})}(X-c)$. Each $\bigcap_{c\in C_{j}(\overline{\mathbb{Q}})}(X-c)$ is a finite intersection of at most $\dim X$ members because of dimension reasons. So $\deg\bigcap_{c\in C_{j}(\overline{\mathbb{Q}})}(X-c)\leq(\deg X)^{\dim X}$ by Bézout’s Theorem. Moreover, each irreducible component of $\bigcap_{c\in C_{j}(\overline{\mathbb{Q}})}(X-c)$ has dimension $\dim X$ because $\dim\mathrm{Stab}(X)=0$. Hence we are in case (ii) by setting $X^{\prime}=\bigcup_{j=1}^{s}\bigcap_{c\in C_{j}}(X-c)$. So we are done. ∎
@@ A.4. Removing $h(X)$
We start by reducing to the case where $X$ generates $A$. Indeed, let $A^{\prime}$ be the abelian subvariety of $A$ generated by XXX-X. Then $X\subseteq A^{\prime}+Q$ for some $Q\in A(\overline{\mathbb{Q}})$. The subgroup $\Gamma^{\prime}$ of $A(\overline{\mathbb{Q}})$ generated by $\Gamma$ and $Q$ has rank $\leq\mathrm{rk}\Gamma+1$. We have $(X-Q)^{\circ}=X^{\circ}-Q$ by definition of the Ueno locus, $(X^{\circ}(\overline{\mathbb{Q}})-Q)\cap\Gamma\subseteq(X^{\circ}(\overline{\mathbb{Q}})-Q)\cap\Gamma^{\prime}=X^{\circ}(\overline{\mathbb{Q}})\cap\Gamma^{\prime}$ and $\deg_{L}(X-Q)=\deg_{L}X$. By Lemma 2.5, (Hyp) yields $\deg_{L}A^{\prime}\leq c^{\prime}(g,\deg_{L}X)$. Therefore if (7.2) holds true for $X-Q\subseteq A^{\prime}$, $L|_{A^{\prime}}$ and $\Gamma^{\prime}\cap A^{\prime}(\overline{\mathbb{Q}})$, then $X^{\circ}(\overline{\mathbb{Q}})\cap\Gamma\leq c(g,\deg_{L}X)^{\mathrm{rk}\Gamma^{\prime}+1}\leq c(g,\deg_{L}X)^{\mathrm{rk}\Gamma+2}$. So we can conclude by replacing $c$ with $c^{2}$. Thus we are reduced to the case where $X$ generates $A$.
From now on, assume $X$ generates $A$. We prove (7.2) by induction on
@@
For an arbitrary $r\geq 1$. Assume (7.2) holds true for $0,1,\ldots,r-1$.
@@
Step 1: Handle large points by both inequalities of Rémond. The goal of this step is to prove the following bound: there exists a constant $c_{6}=c_{6}(g,\deg_{L}X,\deg_{L}A)>0$ such that
(A.5) $\left\{P\in X^{\circ}(\overline{\mathbb{Q}})\cap\Gamma:\hat{h}_{L}(P)\geq c_{5}\max\{1,h(X),h_{\mathrm{Fal}}(A)\}\right\}\leq c_{6}^{\mathrm{rk}\Gamma+1}.$
@@
Consider the subset $\Xi_{j}=\{P_{j+1},\ldots,P_{j+N^{\prime}}\}$ with $j\in\{0,\ldots,N-N^{\prime}\}$; it has $N^{\prime}$ pairwise distinct elements. We claim that $\Xi$ cannot be contained in a proper Zariski closed subset $X^{\prime}\subsetneq X$ with $\deg_{L}X^{\prime}\leq(\deg_{L}X)^{2\dim X}$. Indeed if such an $X^{\prime}$ exists, then by definition of the Ueno locus we have $(X^{\prime})^{\circ}\supseteq X^{\circ}\cap X^{\prime}$. So $\Xi\subseteq(X^{\prime})^{\circ}(\overline{\mathbb{Q}})\cap\Gamma$. As $\dim X^{\prime}<\dim X=r$, we can apply the induction hypothesis (7.2) to each irreducible component of $X^{\prime}$. As $X^{\prime}$ has $\leq(\deg_{L}X)^{2r}$ irreducible components and each component has degree $\leq(\deg_{L}X)^{2r}$, we then get $\Xi\leq(\deg_{L}X)^{2r}c(g,(\deg_{L}X)^{2r},\deg_{L}A)^{\mathrm{rk}\Gamma+1}$. This contradicts our choice of $N^{\prime}$ because $c$ is increasing in all the three variables.
@@
Recall that we have covered $\Gamma\otimes\mathbb{R}$ by at most $\lfloor(1+({8c_{4}})^{1/2})^{\mathrm{rk}\Gamma}\rfloor$ cones and each cone contains $gMN^{\prime}$ points $P\in X^{\circ}(\overline{\mathbb{Q}})\cap\Gamma$ with $\hat{h}_{L}(P)=|P|^{2}\geq c_{5}\max\{1,h(X),h_{\mathrm{Fal}}(A)\}$. Thus
$\#\left\{P\in X^{\circ}(\overline{\mathbb{Q}})\cap\Gamma:\hat{h}_{L}(P)\geq c_{5}\max\{1,h(X),h_{\mathrm{Fal}}(A)\}\right\}$ $\leq(1+({8c_{4}})^{1/2})^{\mathrm{rk}\Gamma}gMN^{\prime}$
@@
Step 2: Remove the dependence on $h(X)$. More precisely, set
@@
(A.10) $\left\{P\in X^{\circ}(\overline{\mathbb{Q}})\cap\Gamma:\hat{h}_{L}(P-P_{0})\geq c_{8}^{2}\max_{1\leq i\leq N^{\prime\prime}}\hat{h}_{L}(P_{i}-P_{0})+c_{9}\max\{1,h_{\mathrm{Fal}}(A)\}\right\}\leq c_{10}^{\mathrm{rk}\Gamma+1}.$
The proof follows closely [Rém00a, Prop. 3.6]. We wish to apply Lemma A.4 to $X-P_{0}$ and the set $S=\{P_{i}-P_{0}:0\leq i\leq N^{\prime\prime}\}$. Let us verify the hypothesis. Let $Y\subseteq X-P_{0}$ with $Y$ equidimensional of dimension $\dim X-1=r-1$ and $S\subseteq Y(\overline{\mathbb{Q}})$, and set $Y^{\prime}:=Y+P_{0}$. Then $P_{i}\in(Y^{\prime})^{\circ}(\overline{\mathbb{Q}})\cap\Gamma$.[5][5][5]As $Y^{\prime}\subseteq X$, we have $(Y^{\prime})^{\circ}\supseteq X^{\circ}\cap Y^{\prime}$ by definition of the Ueno locus. So $P_{i}\in(Y^{\prime}\cap X^{\circ})(\overline{\mathbb{Q}})\cap\Gamma\subseteq(Y^{\prime})^{\circ}(\overline{\mathbb{Q}})\cap\Gamma$. Each irreducible component of $Y^{\prime}$ has dimension $\leq r-1$. Thus we can apply induction hypothesis (7.2) to each irreducible component of $Y^{\prime}$. So $N^{\prime\prime}\leq\sum_{Y^{\prime\prime}}c(g,\deg_{L}Y^{\prime\prime},\deg_{L}A)^{\mathrm{rk}\Gamma+1}$, with $Y^{\prime\prime}$ running over all irreducible components of $Y^{\prime}$. Since $\deg_{L}Y=\deg_{L}Y^{\prime}=\sum_{Y^{\prime\prime}}\deg_{L}Y^{\prime\prime}$ and $c$ is increasing in all the three variables, this bound implies
@@
We claim $\deg_{L}Y>\deg_{L}A(\deg_{L}X)^{2}/g!$. Indeed, assume $\deg_{L}Y\leq\deg_{L}A(\deg_{L}X)^{2}/g!$. Then $N^{\prime\prime}\leq\deg_{L}A(\deg_{L}X)^{2}/g!\cdot c(g,\deg_{L}A(\deg_{L}X)^{2}/g!,\deg_{L}A)^{\mathrm{rk}\Gamma+1}$ because $c$ is increasing in all the three variables. This contradicts the definition of $N^{\prime\prime}$ from (A.9).
@@
Step 3: Prove the following alternative.
@@
    Either $X^{\circ}(\overline{\mathbb{Q}})\cap\Gamma\leq N^{\prime\prime}(8c_{8}+1)^{\mathrm{rk}\Gamma}+c_{10}^{\mathrm{rk}\Gamma+1}$;
@@
    or there exists $Q\in X^{\circ}(\overline{\mathbb{Q}})\cap\Gamma$ such that $\{P\in X^{\circ}(\overline{\mathbb{Q}})\cap\Gamma:\hat{h}_{L}(P-Q)\geq 2c_{9}\max\{1,h_{\mathrm{Fal}}(A)\}\}\leq c_{10}^{\mathrm{rk}\Gamma+1}$.
The proof follows closely [Rém00a, Prop. 3.7]. Assume we are not in case (i), i.e. $X^{\circ}(\overline{\mathbb{Q}})\cap\Gamma>N^{\prime\prime}(8c_{8}+1)^{\mathrm{rk}\Gamma}+c_{10}^{\mathrm{rk}\Gamma+1}$. Let $c_{12}$ be the smallest real number such that there exists $Q\in X^{\circ}(\overline{\mathbb{Q}})\cap\Gamma$ with
(A.11) $\{P\in X^{\circ}(\overline{\mathbb{Q}})\cap\Gamma:|P-Q|\geq c_{12}\}\leq c_{10}^{\mathrm{rk}\Gamma+1}.$
Consider the set $\Xi:=\{P\in X^{\circ}(\overline{\mathbb{Q}})\cap\Gamma:|P-Q|\leq c_{12}\}$ in the $\mathrm{rk}\Gamma$-dimensional Euclidean space $(\Gamma\otimes\mathbb{R},|\cdot|)$. Then $\Xi\geq\#X^{\circ}(\overline{\mathbb{Q}})\cap\Gamma-\#\{P\in X^{\circ}(\overline{\mathbb{Q}})\cap\Gamma:|P-Q|\geq c_{12}\}>N^{\prime\prime}(4c_{8}+1)^{\mathrm{rk}\Gamma}$. By an elementary ball packing argument, $\Xi$ (being a subset of $\Gamma\otimes\mathbb{R}$ contained in a closed ball of radius $c_{12}$ centered at $Q$) is covered by at most $(8c_{8}+1)^{\mathrm{rk}\Gamma}$ balls of radius $c_{12}/4c_{8}$ centered at points in $X^{\circ}(\overline{\mathbb{Q}})\cap\Gamma$; see [Rém00a, Lem. 6.1]. By the Pigeonhole Principle, one of the balls contains $\geq N^{\prime\prime}+1$ points in $X^{\circ}(\overline{\mathbb{Q}})\cap\Gamma$, say $P_{0},P_{1},\ldots,P_{N^{\prime\prime}}$. We have $|P_{i}-P_{0}|\leq c_{12}/2c_{8}$ for each $i$. Thus (A.10) yields
$\left\{P\in X^{\circ}(\overline{\mathbb{Q}})\cap\Gamma:\hat{h}_{L}(P-P_{0})\geq c_{12}^{2}/4+c_{9}\max\{1,h_{\mathrm{Fal}}(A)\}\right\}\leq c_{10}^{\mathrm{rk}\Gamma+1}.$
@@
Step 4: Conclude by the standard packing argument.
Recall our assumption that $X$ generates $A$. The assumption of Theorem 7.1 says that (Hyp pack) holds true for $X$ and $\Gamma$, i.e. we have (7.1).
Assume we are in case (i) from Step 3. Recall the definition of $N^{\prime\prime}=c_{7}^{\mathrm{rk}\Gamma+1}+1$ from (A.9). Then $X^{\circ}(\overline{\mathbb{Q}})\cap\Gamma\leq c^{\mathrm{rk}\Gamma+1}$ for $c:=2\max\{(c_{7}+1)(8c_{8}+1),c_{10}\}$. Hence we can conclude for this case.
@@
(A.12) $\left\{P\in X^{\circ}(\overline{\mathbb{Q}})\cap\Gamma:\hat{h}_{L}(P-Q)\leq 2c_{9}\max\{1,h_{\mathrm{Fal}}(A)\}\right\}\leq c_{0}(1+2\sqrt{2c_{9}c_{0}})^{\mathrm{rk}\Gamma}.$
Thus we have $X^{\circ}(\overline{\mathbb{Q}})\cap\Gamma\leq c_{0}(1+2\sqrt{2c_{9}c_{0}})^{\mathrm{rk}\Gamma}+c_{10}^{\mathrm{rk}\Gamma+1}$. So $X^{\circ}(\overline{\mathbb{Q}})\cap\Gamma\leq c^{\mathrm{rk}\Gamma+1}$ for $c:=2\max\{c_{0},1+2\sqrt{2c_{9}c_{0}},c_{10}\}$. Hence we can conclude are this case.